# Computing Exact Shapley Values in Polynomial Time for Product-Kernel Methods

### v1

@@ 5.2 Explaining distribution discrepancy using MMD with PKeX-Shapley
To illustrate how PKeX-Shapley can explain distributional discrepancies measured by MMD, we conduct two synthetic experiments following the standard two-sample testing setup [10]. Our goal is to attribute the observed MMD between two distributions to individual input variables. We present one of the synthetic experiments below, with the remaining experiments provided in Appendix G.2. In all MMD experiments, we use the RBF kernel with the bandwidth selected via the median heuristic.
We generate datasets $X$ and $Z$, each comprising 20 variables. The first ten variables $X_{1},\dots,X_{10}$ and $Z_{1},\dots,Z_{10}$ are sampled from the same multivariate normal distribution, ensuring identical distributions. The remaining 10 variables $X_{11},\dots,X_{20}$ differ, with variables in $X$ sampled from a multivariate normal distribution and those in $Z$ from a Student’s t-distribution of the same mean. This introduces differences in higher-order moments and covariance structure, resulting in a measurable discrepancy.
Figure 4: KDE plots of Shapley values for the synthetic dataset. Variables 111110101010 reduce MMD via negative contributions, while 1111111120202020 increase it via positive contributions.
@@ 5.3 Explaining HSIC with PKeX-Shapley: Feature selection case study
Lastly, we demonstrate how attributing the HSIC between input features and the target variable to individual features can support feature selection by quantifying their respective contributions to the overall statistical dependence. As a comparison, we also compare our approach with five feature importance methods: HSICLasso [42], Mutual Information (MI), Lasso, K–Best, and Tree Ensemble (with feature permutations). Our experiments were conducted on seven datasets, where we trained Gaussian process (GP) models with an RBF kernel, using only the top 20% of features ranked by each selection method.
@@
6 Conclusion, limitation, and discussion
This work introduces PKeX-Shapley, a polynomial-time algorithm for computing exact Shapley values in product-kernel methods. We show that product kernels naturally induce a functional decomposition, which we exploit to develop a recursive algorithm that avoids the exponential complexity of naive Shapley value computation. This approach enables exact, efficient feature attribution for both predictive models and kernel-based statistical discrepancies, including Maximum Mean Discrepancy (MMD) and Hilbert-Schmidt Independence Criterion (HSIC), providing interpretability of distributional differences and dependence structures. Our method achieves quadratic-time complexity and eliminates the approximation errors inherent in sampling-based estimators (and expectation-based value functions), as demonstrated through experiments on both synthetic and real-world datasets.
While our algorithm reduces computational complexity from exponential to quadratic time, its main limitation is that it applies only to product kernels. This trade-off is largely unavoidable: achieving tractable computation requires imposing structural constraints. As part of future work, we plan to explore whether further relaxation is possible. Another promising direction is to extend our approach to higher-order attribution methods, such as Shapley interaction indices [43, 44]. It is also of interest to investigate connections between our functional decomposition and other explanation techniques, such as partial dependence plots.
Acknowledgement
This project is funded by the Federal Ministry of Education and Research (BMBF). The DAAD Postdoc-NeT-AI short-term research visit scholarship partially supported the first author.
@@ Acknowledgement
  • Newton’s identities : §B presents the main result of Newton’s identities for elementary symmetric polynomials.
  • Proofs: §C provides detailed proofs of key theorems and lemmas used in the main paper.
  • Recursive and numerically stable algorithms for computing Shapley values for product-kernel learning models: §D describes a recursive algorithm for computing Shapley values, as well as an adjusted algorithm for numerical stability and efficiency.
  • Additivity of explanations for learning models, MMD and HSIC: §E discusses the additivity of explanations for the product-kernel methods studied in this paper.
  • Shapley value attribution for HSIC with two multivariate variables §F discusses the Shapley value attribution for HSIC when both random variables are multivariate.
  • Experiments: §G reports experimental setup and extra experiments on explaining product-kernel models.
@@ Appendix A Product kernels in kernel methods
Appendix B Newton’s identities
@@
To explain Newton’s identities, we begin with a specific set of variables $\mathcal{Z}_{4}=\{z_{1},z_{2},z_{3},z_{4}\}$ before generalising it to sets of arbitrary size $\mathcal{Z}_{d}=\{z_{1},z_{2},\ldots,z_{d}\}$. The elementary symmetric polynomials (ESPs) of degree $q$ is defined as
@@ Appendix C Proofs
Theorem 10.
Let the kernel function for feature $j$ be denoted by $k_{j}(x_{j},x^{\prime}_{j})$, and $\prod_{j\in\emptyset}(k_{j}(x_{j},x^{\prime}_{j})-1)=1$ by convention. Then, the following equation holds for product kernel functions:
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{\prime}_{j})=\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{j\in\mathcal{S}}(k_{j}(x_{j},x^{\prime}_{j})-1).$
Proof.
We prove the statement by induction on $|\mathcal{D}|$.
Base case $(|\mathcal{D}|=1)$: Let $\mathcal{D}=\{j\}$. Then $k_{j}(x_{j},x^{\prime}_{j})=1+(k_{j}(x_{j},x^{\prime}_{j})-1),$ as desired.
Inductive step: Assume the equation holds for any set of size $n$. Consider a set $\mathcal{D}$ with size $n+1$, and let $a\in\mathcal{D}$. We write:
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{\prime}_{j})=k_{a}(x_{a},x^{\prime}_{a})\prod_{j\in\mathcal{D}\setminus\{a\}}k_{j}(x_{j},x^{\prime}_{j}).$
Using the induction hypothesis for $\mathcal{D}\setminus\{a\}$, we have
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{\prime}_{j})=k_{a}(x_{a},x^{\prime}_{a})\sum_{\mathcal{T}\subseteq\mathcal{D}\setminus\{a\}}\prod_{j\in\mathcal{T}}(k_{j}(x_{j},x^{\prime}_{j})-1).$
Since $k_{a}(x_{a},x^{\prime}_{a})=(k_{a}(x_{a},x^{\prime}_{a})-1)+1$, expanding gives:
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{\prime}_{j})=\left[(k_{a}(x_{a},x^{\prime}_{a})-1)+1\right]\sum_{\mathcal{T}\subseteq\mathcal{D}\setminus\{a\}}\prod_{j\in\mathcal{T}}(k_{j}(x_{j},x^{\prime}_{j})-1).$
Expanding this product, we obtain:
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{\prime}_{j})=\sum_{\mathcal{T}\subseteq\mathcal{D}\setminus\{a\}}(k_{a}(x_{a},x^{\prime}_{a})-1)\prod_{j\in\mathcal{T}}(k_{j}(x_{j},x^{\prime}_{j})-1)\>+\sum_{\mathcal{T}\subseteq\mathcal{D}\setminus\{a\}}\prod_{j\in\mathcal{T}}(k_{j}(x_{j},x^{\prime}_{j})-1).$
The first sum on the right-hand side of the equation covers all subsets that contain $a$, and the second sum covers all subsets that do not contain $a$. Thus, together they sum over all subsets $\mathcal{\mathcal{S}}\subseteq\mathcal{\mathcal{D}}$:
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{\prime}_{j})=\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{j\in\mathcal{S}}(k_{j}(x_{j},x^{\prime}_{j})-1).$
Thus, by induction, the theorem is proved. ∎
We now provide the proofs for all the theorems and lemmas in the manuscript.
C.1 Proof of Theorem 1
Using the product kernel and its property in equation (2), the decision function can be rewritten as:
$f(\boldsymbol{x})=\sum_{i=1}^{n}\alpha_{i}\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{(i)}_{j}).$
Expanding the product using Theorem 10:
$\prod_{j\in\mathcal{D}}k_{j}(x_{j},x^{(i)}_{j})=\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}.$
Substituting this expansion into the decision function:
$f(\boldsymbol{x})=\sum_{i=1}^{n}\alpha_{i}\sum_{\mathcal{S}\subseteq\mathcal{D}}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}.$
Rearranging the summations:
$f(\boldsymbol{x})=\sum_{\mathcal{S}\subseteq\mathcal{D}}\sum_{i=1}^{n}\alpha_{i}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}.$
This establishes the functional decomposition:
$f(\boldsymbol{x})=\sum_{\mathcal{S}\subseteq\mathcal{D}}f_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}}),$
where each term is uniquely defined as:
$f_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}})=\sum_{i=1}^{n}\alpha_{i}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}.$
Since each subset $\mathcal{S}$ contributes independently to the sum and there is no overlap between terms, this decomposition is exact and uniquely determined by the structure of the product kernel. ∎
C.2 Proof of Proposition 3
Based on Definition 2, we have:
$v_{\boldsymbol{x}}(\mathcal{S})=\sum_{\mathcal{T}\subseteq\mathcal{S}}f_{\mathcal{T}}(\boldsymbol{x}_{\mathcal{T}}),$
where $v_{\boldsymbol{x}}(\mathcal{S})$ is a unique value function that quantifies the contribution of feature set $\mathcal{S}$ to the prediction. Substituting $f_{\mathcal{T}}$ from the decomposition, one gets:
$v_{\boldsymbol{x}}(\mathcal{S})=\sum_{\mathcal{T}\subseteq\mathcal{S}}\sum_{i=1}^{n}\alpha_{i}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}=\sum_{i=1}^{n}\alpha_{i}\sum_{\mathcal{T}\subseteq\mathcal{S}}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}.$
Using Theorem 10,
$\sum_{\mathcal{T}\subseteq\mathcal{S}}\prod_{j\in\mathcal{T}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}=\prod_{j\in\mathcal{S}}k_{j}(x_{j},x^{(i)}_{j}),$
which simplifies the value function to
$v_{\boldsymbol{x}}(\mathcal{S})=\sum_{i=1}^{n}\alpha_{i}\prod_{j\in S}k_{j}(x_{j},x^{(i)}_{j})=\sum_{i=1}^{n}\alpha_{i}k_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}^{(i)}_{\mathcal{S}}),$
and that completes the proof. ∎
C.3 Proof of Theorem 4

### v2

@@ 5.2 Explaining distribution discrepancy using MMD with PKeX-Shapley
To illustrate how PKeX-Shapley can explain distributional discrepancies measured by MMD, we conduct two synthetic experiments following the standard two-sample testing setup (Schrab, 2025). Our goal is to attribute the observed MMD between two distributions to individual input variables. We present one of the synthetic experiments below, with the remaining experiments provided in Appendix H.3. In all MMD experiments, we use the RBF kernel with the bandwidth selected via the median heuristic method. We generate datasets $X$ and $Z$, each comprising 20 variables. The first ten variables $X_{1},\dots,X_{10}$ and $Z_{1},\dots,Z_{10}$ are sampled from the same multivariate normal distribution, ensuring identical distributions. The remaining 10 variables $X_{11},\dots,X_{20}$ differ, with variables in $X$ sampled from a multivariate normal distribution and those in $Z$ from a Student’s t-distribution of the same mean. This introduces differences in higher-order moments and covariance structure, resulting in a measurable discrepancy.
Figure 5: KDE plots of Shapley values for the synthetic dataset. Variables 111010 reduce MMD via negative contributions, while 11112020 increase it via positive contributions.
@@ 5.3 Explaining HSIC with PKeX-Shapley: Feature selection case study
Lastly, we demonstrate how attributing the HSIC between input features and the target variable to individual features can support feature selection by quantifying their respective contributions to the overall statistical dependence. As a comparison, we also compare our approach with five feature importance methods: HSICLasso (Climente-González et al., 2019), Mutual Information (MI), Lasso, K–Best, and Tree Ensemble (with feature permutations). Our experiments were conducted on seven datasets, where we trained Gaussian process (GP) models with an RBF kernel, using only the top 20% of features ranked by each selection method.
@@
  • Newton’s identities : §C presents the main result of Newton’s identities for elementary symmetric polynomials.
  • Proofs: §D provides detailed proofs of key theorems and lemmas used in the main paper.
  • Recursive and numerically stable algorithms for computing Shapley values for product-kernel learning models: §E describes a recursive algorithm for computing Shapley values, as well as an adjusted algorithm for numerical stability and efficiency.
  • Additivity of explanations for learning models, MMD and HSIC: §F discusses the additivity of explanations for the product-kernel methods studied in this paper.
  • Shapley value attribution for HSIC with two multivariate variables §G discusses the Shapley value attribution for HSIC when both random variables are multivariate.
  • Experiments: §H reports experimental setup and extra experiments on explaining product-kernel models.
@@ Appendix A Product kernels in kernel methods
Appendix B Value Functions for Product Kernels: Baseline and Interventional
We now discuss the interventional and baseline value functions for product-kernel methods and their relevance and affinity with the proposed value function. For a product-kernel model, the decision function can be expressed as
$f(\boldsymbol{x})=\boldsymbol{\alpha}^{\top}\big(k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}})\odot k_{\mathcal{S}^{c}}(\mathbf{X}_{\mathcal{S}^{c}},\boldsymbol{x}_{\mathcal{S}^{c}})\big),$
where $\odot$ denotes the Hadamard product, $k_{\mathcal{S}}$ is the product kernel restricted to subset $\mathcal{S}$, and $\mathcal{S}^{c}$ is its complement.
B.1 Baseline Value Function
The baseline value function removes the effect of features outside the coalition $\mathcal{S}$ by replacing them with fixed baseline values. Formally, let $\boldsymbol{x}_{S^{c}}^{b}$ denote the baseline values for the excluded features. Then the baseline value function is defined as
$v_{\boldsymbol{x}}^{\text{base}}(\mathcal{S})$ $=f([\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}_{S^{c}}^{b}])$
$=\boldsymbol{\alpha}^{\top}\mathbf{K}([\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}_{S^{c}}^{b}],\mathbf{X})=\boldsymbol{\alpha}^{\top}\big(k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}})\odot k_{\mathcal{S}^{c}}(\mathbf{X}_{\mathcal{S}^{c}},\boldsymbol{x}^{b}_{\mathcal{S}^{c}})\big).$
In model-agnostic settings, the choice of baseline values is often application-dependent (e.g., mean input, zero vector, or reference sample). For product-kernel methods, however, we have an explicit multiplicative structure. Instead of fixing raw feature values, we can directly mask the kernel factors corresponding to excluded features. Specifically, setting a kernel factor to one corresponds to the least informative contribution, since it makes the kernel computation independent of that feature. Thus, the baseline operation in product kernels amounts to replacing $k_{j}(x_{j},x_{j}^{(i)})$ with $1$ for all $j\in S^{c}$. This yields
$v_{\boldsymbol{x}}(\mathcal{S})$ $=\boldsymbol{\alpha}^{\top}k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}}),$
which is exactly the functional baseline value function.
In words, rather than choosing baseline feature values in input space, product-kernel methods define the baseline at the level of the kernel: excluded features are masked by setting their kernel contributions to $1$. This ensures that only features in $\mathcal{S}$ influence the similarity measure.
B.2 Interventional Value Function
An alternative approach is the interventional value function, which removes the effect of features outside $\mathcal{S}$ by replacing them with their average contribution under the data distribution. Formally, it is defined as
$v_{\boldsymbol{x}}^{\text{int}}(\mathcal{S})$ $=\mathbb{E}_{\boldsymbol{x}_{\mathcal{S}^{c}}}\left[f(\boldsymbol{x})\mid\boldsymbol{x}_{\mathcal{S}}\right]$
$=\mathbb{E}_{\boldsymbol{x}_{\mathcal{S}^{c}}}\left[\boldsymbol{\alpha}^{\top}k([\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}^{c}}],\mathbf{X})\right]$
$=\boldsymbol{\alpha}^{\top}\,\mathbb{E}_{\boldsymbol{x}_{\mathcal{S}^{c}}}\left[k([\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}^{c}}],\mathbf{X})\right].$
Exploiting the product structure of the kernel, this simplifies to
$v_{\boldsymbol{x}}^{\text{int}}(\mathcal{S})=\boldsymbol{\alpha}^{\top}\Big(k_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}},\mathbf{X}_{\mathcal{S}})\odot\mathbb{E}_{\boldsymbol{x}_{\mathcal{S}^{c}}}\big[k_{\mathcal{S}^{c}}(\boldsymbol{x}_{\mathcal{S}^{c}},\mathbf{X}_{\mathcal{S}^{c}})\big]\Big).$ (5)
To summarize, the interventional value function removes a feature’s contribution by averaging it out, while our baseline-inspired value function directly eliminates the feature’s influence by setting its kernel component to one.
This highlights the conceptual and practical difference: the interventional function is defined through expectations with respect to the data distribution, whereas our proposed value function exploits the explicit multiplicative decomposition of product kernels and provides an efficient, distribution-free alternative.
B.3 Value Function as a Projection
We now present the value function as the orthogonal projection of the predictor $f\in\mathcal{H}$ onto the closed subspace $\mathcal{H}_{\mathcal{S}}$ of $\mathcal{S}$-only functions. That is, for a coalition $\mathcal{S}\subseteq\mathcal{D}$, the worth of $\mathcal{S}$ for an input $\boldsymbol{x}$ is defined by
$v_{\boldsymbol{x}}(\mathcal{S}):=(\Pi_{\mathcal{S}}f)(\boldsymbol{x}),$
where $\Pi_{\mathcal{S}}$ denotes the orthogonal projection operator from $\mathcal{H}$ to $\mathcal{H}_{\mathcal{S}}$.
Theorem 9 (Orthogonal projection onto the $\mathcal{S}$-only subspace).
Let $(\mathcal{H},k)$ be an RKHS of real-valued functions on $X=X_{1}\times\cdots\times X_{d}$. For a fixed index set $\mathcal{S}\subseteq\{1,\dots,d\}$, define the $\mathcal{S}$-only subspace
$\mathcal{H}_{\mathcal{S}}\;:=\;\{\,g\in\mathcal{H}:\ g(\boldsymbol{x})=g(\boldsymbol{x}^{\prime})\ \text{whenever}\ \boldsymbol{x}_{\mathcal{S}}=\boldsymbol{x}^{\prime}_{\mathcal{S}}\,\}.$
Then:
  1. 1.
    $\mathcal{H}_{\mathcal{S}}$ is a closed linear subspace of $\mathcal{H}$. Hence, the orthogonal projection $\Pi_{\mathcal{S}}:\mathcal{H}\to\mathcal{H}_{\mathcal{S}}$ exists.
  2. 2.
    Define the kernel
    $k_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}^{\prime}_{\mathcal{S}})\ :=\ \big\langle\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x}),\ \Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x}^{\prime})\big\rangle_{\mathcal{H}}.$
    Then $k_{\mathcal{S}}$ is positive definite on $X_{\mathcal{S}}$, and for every $\boldsymbol{x}\in X$,
    $\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x})\;=\;k_{\mathcal{S}}(\cdot,\boldsymbol{x}_{\mathcal{S}})\ \in\mathcal{H}_{\mathcal{S}}.$
  3. 3.
    For any $f=\sum_{i=1}^{n}\alpha_{i}\,k(\cdot,\boldsymbol{x}^{(i)})\in\mathcal{H}$,
    $\Pi_{\mathcal{S}}f\;=\;\sum_{i=1}^{n}\alpha_{i}\,k_{\mathcal{S}}(\cdot,\boldsymbol{x}^{(i)}_{\mathcal{S}}),\qquad(\Pi_{\mathcal{S}}f)(\boldsymbol{x})\;=\;\sum_{i=1}^{n}\alpha_{i}\,k_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}^{(i)}_{\mathcal{S}})$.
Proof 1.
(a) Linearity of $\mathcal{H}_{\mathcal{S}}$ is immediate. To prove closedness, let $g_{n}\in\mathcal{H}_{\mathcal{S}}$ with $g_{n}\to g$ in $\mathcal{H}$. Since evaluation functionals are continuous in an RKHS,
$|g_{n}(\boldsymbol{x})-g(\boldsymbol{x})|\;\leq\;\|g_{n}-g\|_{\mathcal{H}}\,\sqrt{k(\boldsymbol{x},\boldsymbol{x})}\ \to\ 0.$
If $\boldsymbol{x}_{\mathcal{S}}=\boldsymbol{x}^{\prime}_{\mathcal{S}}$, then $g_{n}(\boldsymbol{x})=g_{n}(\boldsymbol{x}^{\prime})$ for all $n$, hence $g(\boldsymbol{x})=g(\boldsymbol{x}^{\prime})$. Thus $g\in\mathcal{H}_{\mathcal{S}}$, proving closedness.
(b) For fixed $\boldsymbol{x}$, define the functional on $\mathcal{H}_{\mathcal{S}}$, $\Lambda_{\boldsymbol{x}}(g):=g(\boldsymbol{x})=g(\boldsymbol{x}_{\mathcal{S}})$. It is bounded, so by the Riesz representation there is a unique $p_{\boldsymbol{x}}\in\mathcal{H}_{\mathcal{S}}$ such that
$g(\boldsymbol{x})=\langle g,p_{\boldsymbol{x}}\rangle_{\mathcal{H}},\qquad\forall g\in\mathcal{H}_{\mathcal{S}}.$
But also $\langle g,\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x})\rangle_{\mathcal{H}}=\langle g,k(\cdot,\boldsymbol{x})\rangle_{\mathcal{H}}=g(\boldsymbol{x})$. Thus $p_{\boldsymbol{x}}=\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x})$, and we identify
$k_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}},\boldsymbol{x}^{\prime}_{\mathcal{S}})=\langle p_{\boldsymbol{x}},p_{\boldsymbol{x}^{\prime}}\rangle_{\mathcal{H}}=\langle\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x}),\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x}^{\prime})\rangle_{\mathcal{H}}.$
Therefore $\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x})=k_{\mathcal{S}}(\cdot,\boldsymbol{x}_{\mathcal{S}})$.
(c) By linearity of $\Pi_{\mathcal{S}}$,
$\Pi_{\mathcal{S}}f=\sum_{i=1}^{n}\alpha_{i}\,\Pi_{\mathcal{S}}k(\cdot,\boldsymbol{x}^{(i)})=\sum_{i=1}^{n}\alpha_{i}\,k_{\mathcal{S}}(\cdot,\boldsymbol{x}^{(i)}_{\mathcal{S}})$.
Evaluating at $\boldsymbol{x}$ gives the desired expansion.
This definition is meaningful because the projection $\Pi_{\mathcal{S}}f$ is the unique element of $\mathcal{H}_{\mathcal{S}}$ that is closest to $f$ in the RKHS norm. Hence, $(\Pi_{\mathcal{S}}f)(\boldsymbol{x})$ represents the prediction of the best possible $\mathcal{S}$-only approximation of $f$, measured in the same geometry that the model itself inhabits.
Appendix C Newton’s identities
@@
To explain Newton’s identities, we begin with a specific set of variables $\mathcal{Z}_{4}=\{z_{1},z_{2},z_{3},z_{4}\}$ before generalizing it to sets of arbitrary size $\mathcal{Z}_{d}=\{z_{1},z_{2},\ldots,z_{d}\}$. The elementary symmetric polynomials (ESPs) of degree $q$ is defined as
@@ Appendix D Proofs
D.1 Proof of Theorem 3